find the radius of a circle with an area of 80cm 2
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

find the radius of a circle with an area of 80cm 2
pi*r^2 = 80
r^2 = 80/pi
r = 4sqrt(3)/pi
---------------
Cheers,
Stan H.
-------------

Answer by addingup(3677)

What is the area? It is:
r^2*Pi = area
now we work it backwards:
r^2*Pi = 80
r^2 = 80/Pi
r^2 = 25.46
sqrt(r^2) = sqrt(25.46)
r = 5.05